OpenStudy (broskishelleh):

Determine the equation of the line, in standard form, that will get your spacecraft from the Launch Area to Point A. Launch Area:___(1, 2)___ Point A:___(0, 3)___

OpenStudy (broskishelleh):

@goformit100

OpenStudy (anonymous):

x+y=3 thats the answer I already did this assignment. :)

OpenStudy (broskishelleh):

Are you sure? @Morgan27 What are the steps?

OpenStudy (broskishelleh):

May someone please explain how to complete this? Please...

OpenStudy (broskishelleh):

@undeadknight26 Mah homie, can ya explain?

undeadknight26 (undeadknight26):

Is im lazy an excuse? lololol jk broda ill see what i can whip up for ya

OpenStudy (broskishelleh):

I'm lazy is an excuse used to procastinate hours and hours of accumulated assessments to a later moment and/or time.

OpenStudy (broskishelleh):

Wow my openstudy is lagging

undeadknight26 (undeadknight26):

Touche =.= Explanation: @Morgan27 is right and im procrastinating on this question because im behind in spanish -_- lol ask @hugsnkisses he did this

undeadknight26 (undeadknight26):

Wrong person @hugsntoughs

undeadknight26 (undeadknight26):

Awe crackers he logged...get him on gmail.

OpenStudy (broskishelleh):

I know I did

OpenStudy (anonymous):

well for question number one on this assignment it asked Graph the coordinates of the specifics points in space your spacecraft will travel to. right? ok so I graphed the points and for #1 I used the slope form for the (1,2) and (0,3) to find the slope…. and then to find the y intercept used coordinate (0,3) and the slope in the formula y=mx+b to finally get y=-x+3

undeadknight26 (undeadknight26):

@hugsnotughs

OpenStudy (anonymous):

but they wanted it in standard form you would add x to y and get x+y=3

OpenStudy (hugsnotughs):

Thanks for tagging me Mr. Grey. So far @Morgan27 is doing an excellent job in explaining.

undeadknight26 (undeadknight26):

@hugsnotughs @hugsnotughs @hugsnotughs@hugsnotughs@hugsnotughs@hugsnotughs@hugsnotughs

undeadknight26 (undeadknight26):

@hugsnotughs

OpenStudy (dbzfan836):

I agree

OpenStudy (hugsnotughs):

Really Mr. Grey...that was unnecessary.

OpenStudy (anonymous):

Hope that helped a little lol

undeadknight26 (undeadknight26):

English is boring...just reading about love -_- And making me destroy pilows =.=

OpenStudy (dbzfan836):

Lol rly Mr. Grey?

undeadknight26 (undeadknight26):

I am Mr. Grey ;D You can call me mista sexy if you would like.

OpenStudy (dbzfan836):

You really should remove that comment

OpenStudy (hugsnotughs):

Yes, Tariq's nickname is Mr. Grey. And Mr. Grey, you need to find the emotion of the poem. You aren't reading that all day long, learn how to focus and get off of OS for a while. ;)

OpenStudy (dbzfan836):

I know that

OpenStudy (hugsnotughs):

Lol, poor thing, we're spamming @Broskishelleh 's post. ;-;

OpenStudy (broskishelleh):

It's fine, @Morgan27 did such a great job

OpenStudy (broskishelleh):

Have fun

undeadknight26 (undeadknight26):

I know but love sickens me =.=

OpenStudy (hugsnotughs):

Mr. Grey, soon you'll be in love with someone like me, so not to worry. Love comes in the strangest ways and int he strangest times.

OpenStudy (anonymous):

Thank you :)